NARCC H240) NMR (Cq HOO)

Use the spectra below to fill in the boxes for the reaction above.

If it isn't clear box J is (C11 H24 O) and box I is (C9 H10 O)

Molecule 9H Зн 2н 2н PPMMolecule ) 9H ЗН ін 2H 2H Е -
Answer #1

84 -4-2+1 2 -12-0 S3.4 CSI3u S2-10mu) CNTIbTS S1.4C9728) CqHioO 9-(041 = l0-5 5441 ca Higo CVBE ~ 9-194 10-9=4 . 25C tI2U) S1